How many grams of solid potassium fluoride should be added to 0.500 L of a 0.284 M hydrofluoric acid solution to prepare a buffer with a pH of 2.507 ?  
Answer #1

pKa of HF = 3.2 (Literature value)

mole of HF = 0.500 L * 0.284 M = 0.142 mole.

solution is buffer solution.

Using Henderson equation is

pH = pKa + log [salt] / [acid]

or

2.507 = 3.2 + log [salt] / 0.142

or

[salt] = 0.0288 mole

thus

grams of solid potassium fluoride added = mole * molar mass = 0.0288 mole * 58.0967 g/mole = 1.67 g (answer)

If you get the pKa of HF then you also get the accurate result.